# CDP: Towards Robust Autoregressive Visuomotor Policy Learning via Causal Diffusion

### v1

@@ Abstract
Diffusion Policy (DP) enables robots to learn complex behaviors by imitating expert demonstrations through action diffusion. However, in practical applications, hardware limitations often degrade data quality, while real-time constraints restrict model inference to instantaneous state and scene observations. These limitations seriously reduce the efficacy of learning from expert demonstrations, resulting in failures in object localization, grasp planning, and long-horizon task execution. To address these challenges, we propose Causal Diffusion Policy (CDP), a novel transformer-based diffusion model that enhances action prediction by conditioning on historical action sequences, thereby enabling more coherent and context-aware visuomotor policy learning. To further mitigate the computational cost associated with autoregressive inference, a caching mechanism is also introduced to store attention key-value pairs from previous timesteps, substantially reducing redundant computations during execution. Extensive experiments in both simulated and real-world environments, spanning diverse 2D and 3D manipulation tasks, demonstrate that CDP uniquely leverages historical action sequences to achieve significantly higher accuracy than existing methods. Moreover, even when faced with degraded input observation quality, CDP maintains remarkable precision by reasoning through temporal continuity, which highlights its practical robustness for robotic control under realistic, imperfect conditions.
@@ 1 Introduction
    To enable efficient real-time inference, we introduce a caching mechanism that stores and reuses attention key-value pairs from prior timesteps, substantially reducing computational overhead in autoregressive action generation.
@@ 3.1 Causal Action Generation
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation along the temporal axis and $D_{\theta}$ denotes the whole denoising process, based on our Causal Action Generation Module with the parameter of $\theta$.
@@
where $\tilde{\textbf{A}}^{\texttt{perturb}}$ denotes the perturbed historical actions and $\textbf{N}_{\sigma}\sim\mathcal{N}(0,\sigma^{2})$ denotes the small-scale noise sequence with the variance of $10<\sigma<$ 1. By introducing $\textbf{N}_{\sigma}$, we create a more robust training environment that simulates the occasions where Historical Actions $\tilde{\textbf{A}}$ may be imperfect or noisy, forcing the model to focus on the coarse-grained temporal dynamic context from the Historical Actions $\tilde{\textbf{A}}$, instead of depending on their concrete values. These coarse-grained temporal dynamics serve as a complement to the degraded observations O, guiding the subsequent denoising process.
Causal Action Generation Module. As depicted in Fig. 2 (a), this module is composed of $P$ blocks. Within each block, we first inject the temporal dynamic context from the Historical Actions $\tilde{\textbf{A}}$ into the Denoising Targets N through Causal Temporal Attention (CTA), generating intermediate features $\mathbf{N}_{t}$ Following this, Visual-Action Cross Attention (VACA) is utilized to incorporate spatial constraints from the degraded observation O into these $\mathbf{N}_{t}$, generating intermediate features $\mathbf{N}_{ts}$. These $\mathbf{N}_{ts}$ are then further processed by a Multi-layer Perceptron (MLP). It is worth noting that we place the denoising timestep embedding $t$ in the VACA layer, rather than the CTA layer in order to ensure the cached features of the Historical Actions $\tilde{\textbf{A}}$ can be correctly shared during the whole denoising process (Sec. 3.2). The entire process within each block can be described as follows,
@@
where $\mathbf{N}_{o}$ represents the output features of this block. LN denotes the Layer Normalization operation, and Enc denotes the 2D or 3D data encoder, which mainly used to extract the information from 2D or 3D observation.
@@ 3.2 Chunk-wise Autoregressive Inference with Cache Sharing
$\mathbf{A}^{k}\sim p_{\theta}(\mathbf{A}^{k}|\mathbf{N}^{k},\tilde{\textbf{A}}^{k}),$ (6)
@@
Cache Sharing. In the AR-k step, the Historical Actions $\tilde{\textbf{A}}^{k}$ of length $L$ can be divided into two parts: Cached Historical Actions $\tilde{\textbf{A}}^{k}_{0:l-1}$ and Uncached Historical Actions $\tilde{\textbf{A}}^{k}_{l:L-1}$, based on whether their Key and Value representations have been cached up to this step. Here, $l$ denotes the length of the Cached Historical Actions. During denoising, the Key and Value representations for the Cached Historical Actions $\tilde{\textbf{A}}^{k}_{0:l-1}$ have been precomputed in prior autoregressive steps and can be directly used in attention computation. We denote them as $\mathbf{K}_{0:l-1}^{(0)}$ and $\mathbf{V}_{0:l-1}^{(0)}$, where ()($\cdot$)( ⋅ ) indicates the denoising timestep and $(0)$ means these Key and Value representations correspond to the final Target Actions generated in prior steps. For the Uncached Historical Actions $\tilde{\textbf{A}}^{k}_{l:L-1}$ and Denoising Targets $\textbf{N}^{k}$, we utilize a QKV_MLP layer to extract their Query, Key, and Value representations,
[…]
@@ 4.1 Experiment Setup
Evaluation Metric. Following the established metrics [27], each experiment was conducted across three independent trials, utilizing seed values of 0, 1, and 2, in the Adroit, DexArt, and MetaWorld benchmarks. For each seed, the policy was evaluated over 20 episodes every 200 training epochs, and the mean of the top 5 success rates was computed. The final reported performance consists of the mean and standard deviation of these success rates across above three seeds. In the Robofactory benchmark, each experiment was executed using only one seed value, i.e. 0, and was evaluated over 100 episodes at the 300th training epoch. This comprehensive evaluation metrics ensures a fair comparison between our CDP and baseline methods, providing a clear assessment of the performance improvements achieved by our approach.
4.2 Result Analysis
@@ 4.3 Ablation Study
Figure 4: Ablation study about the quality of observation data. CDP consistently outperforms DP across all experimental conditions.
Ablation Study on the Quality of Observation Data. To thoroughly examine the robustness of our proposed method under different levels of observation quality, we conduct experiments on the Lift Barrier task of the Robofactory benchmark using point clouds of varying resolutions, specifically with 32, 64 and 128 points. We then compare the success rates of our method with those of DP3 across these point clouds. Fig. 4 reveals that our method demonstrates significantly superior robustness to degraded observations. As the resolution of the point clouds decreases, the success rate of DP3 drops markedly, whereas our proposed method maintains a consistently high success rate. This resilience can be attributed to the model’s ability to effectively utilize historical action sequences, thereby compensating for the lack of spatial constraints from degraded observations.
@@ 5 Real-World Experiments
5.1 Experiment Setup
Workspace. As illustrated in Fig. 6, our real-world experiments were conducted using a RealMan robotic arm fitted with a DH AG95 gripper. A stationary top-down Intel RealSense D435i RGB-D camera was employed to provide a global RGB image of the workspace. All hardware components were connected to a workstation equipped with an NVIDIA 3090 Ti GPU. This workspace enabled efficient data acquisition and evaluation of the robotic system’s performance.
Demonstrations. The demonstrations utilized in our experiments were meticulously gathered via human teleoperation, facilitated by advanced vision-based retargeting techniques. For each task, we collected a total of 50 demonstrations, each carefully selected to encapsulate the fundamental skills and critical interactions necessary for achieving successful task outcomes. This curated approach not only ensures that our dataset remains manageable in size but also guarantees its representativeness of the inherent complexities and challenges associated with the tasks.
5.2 Results Analysis
@@ 6 Conclusion
In conclusion, our proposed CDP effectively addresses key challenges in imitation learning for robotic manipulation by leveraging a causal transformer policy that incorporates Historical Actions. This design enables the model to capture essential temporal dependencies and contextual cues, leading to more coherent and adaptive behavior. Through comprehensive experiments, CDP has demonstrated superior performance over existing methods in both simulated and real-world tasks, showcasing its robustness and effectiveness. Future work will explore further enhancements to handle even more complex environments and integrate advanced techniques to improve generalization and performance.
Limitation. This work experimentally validated the effectiveness of causal action generation under conditions of degraded data quality. However, several limitations remain to be addressed. Firstly, while our approach demonstrates significant improvements in action prediction by leveraging historical actions, an efficient fusion strategy of Historical Actions with instantaneous observations remains a challenging problem. Specifically, developing a more effective mechanism to achieve seamless and computationally efficient complementarity between Historical Actions and current observations is still an open question. Secondly, this work has not explored tasks with extremely long horizons. The application of our method to such tasks, which require sustained and coherent action planning over extended periods, is left for future investigation.
@@ Appendix A Simulation Experiments Details
A.1 Settings
In this part, we provide detailed descriptions of all settings in our simulation experiments, including those related to the training strategy, as well as the inputs (Observations and Historical Actions) and outputs (Target Actions).
Table 3: All Settings in Simulation Experiments. The notation $\mathcal{N}(0,1)$ denotes Gaussian noise images characterized by a mean of 0 and a variance of 1. To ensure action coherence throughout the inference process, we utilize only the initial segment of the predicted target actions as the final target actions, termed the Valid Parts, while the subsequent segment is discarded, referred to as the Redundant Parts.
@@ A.1 Settings
Number of Point Clouds 512
@@
Specifically, we made the following changes in the simulated experiments conducted on the RoboFactory benchmark,
  • In the 2D scenario, images were resized to (64, 64, 3) instead of (84, 84, 3). And in the 3D scenario, the number of point clouds was reduced to 128, deviating from the standard configuration.
@@
  • The training epoch is set to 300 in RoboFactory benchmark, instead of 3000.
A.2 Supplementary Experiments on Ablation Study
In this section, we conduct experiments on the Lift Barrier task of the Robofactory benchmark to investigate the impact of Historical Length and Target Length on model performance.
Ablation Study on the Length of Historical Actions. Table 4 presents the success rates of the model for different configurations of Historical Actions length. From this table, we find that a longer sequence of Historical Actions tends to be beneficial in enhancing model performance.
Table 4: Quantitative results of our proposed methods across different length of Historical Actions. The format “? - $*_{1}$ $*_{2}$” is used to denote the following: ? represents the length of Historical Actions. $*_{1}$ and $*_{2}$ represent the length of valid and redundant parts of Target Actions, respectively.
Method ? - 4 - 4 ? - 4 - 8
4 8 16 4 8 16
CDP3 0% 0% 36% 0% 6% 37%
CDP 0% 0% 37% 0 0% 39%
Ablation Study on the Length of Target Actions. Table 5 illustrates that longer Target Actions will provide higher gains for the model. When combined with the analysis in Table 4, we observe that the optimal length of Target Actions should be determined in conjunction with the length of Historical Actions. Only well-matched combinations of these action lengths can effectively enhance model performance.
Table 5: Quantitative results of our proposed method on different length of Target Actions. The format “8 - ? - 8” is used to denote the following: ? represents the length of the valid parts of Target Actions, while both the length of Historical Actions and the length of the redundant parts of Target Actions are set to 8.
Method 8 - ? - 8
2 4 8
CDP3 0% 6% 33%
CDP 0% 0% 38%
A.3 Qualitative Results of Simulation Experiments
@@
Figure 5: Visualization of real-world and simulated experimental results, including Stacking Cubes, Collecting Objects, Push T, Lift Barrier, Dexart Laptop, Metaworld Box Close and Adroit Door task from the top to the bottom.
@@ B.1 Workspace
We illustrate our real-world experimental workspace as Fig. 6, which mainly includes an RGB camera and a robotic arm.
@@
B.2 Settings
In this part, we list all settings in our real-world experiments in Table 6. Similar with the simulated experiments, these settings involve the training strategy, as well as the inputs (Observations and Historical Actions) and outputs (Target Actions).
Table 6: All Settings in Real-World Experiments

### v2

@@ Abstract
Diffusion Policy (DP) enables robots to learn complex behaviors by imitating expert demonstrations through action diffusion. However, in practical applications, hardware limitations often degrade data quality, while real-time constraints restrict model inference to instantaneous state and scene observations. These limitations seriously reduce the efficacy of learning from expert demonstrations, resulting in failures in object localization, grasp planning, and long-horizon task execution. To address these challenges, we propose Causal Diffusion Policy (CDP), a novel transformer-based diffusion model that enhances action prediction by conditioning on historical action sequences, thereby enabling more coherent and context-aware visuomotor policy learning. To further mitigate the computational cost associated with autoregressive inference, a caching mechanism is also introduced to store attention key-value pairs from previous timesteps, substantially reducing redundant computations during execution. Extensive experiments in both simulated and real-world environments, spanning diverse 2D and 3D manipulation tasks, demonstrate that CDP uniquely leverages historical action sequences to achieve significantly higher accuracy than existing methods. Moreover, even when faced with degraded input observation quality, CDP maintains remarkable precision by reasoning through temporal continuity, which highlights its practical robustness for robotic control under realistic, imperfect conditions. Project page: https://gaavama.github.io/CDP/
@@ 1 Introduction
    To enable efficient real-time inference, we introduce a cache sharing mechanism that stores and reuses attention key-value pairs from prior timesteps, substantially reducing computational overhead in autoregressive action generation.
@@ 3.1 Causal Action Generation
where $\cdot$$\cdot$] denotes the concatenation operation along the temporal axis and $D_{\theta}$ denotes the whole denoising process, based on our Causal Action Generation Module with the parameter of $\theta$.
@@
where $\tilde{\textbf{A}}^{\texttt{perturb}}$ denotes the perturbed historical actions and $\textbf{N}_{\sigma}\sim\mathcal{N}(0,\sigma^{2})$ denotes the small-scale noise sequence with the variance of $0<\sigma<1$. By introducing $\textbf{N}_{\sigma}$, we create a more robust training environment that simulates the occasions where Historical Actions $\tilde{\textbf{A}}$ may be imperfect or noisy, forcing the model to focus on the coarse-grained temporal dynamic context from the Historical Actions $\tilde{\textbf{A}}$, instead of depending on their concrete values. These coarse-grained temporal dynamics serve as a complement to the degraded observations O, guiding the subsequent denoising process.
Causal Action Generation Module. As depicted in Fig. 4 (a), this module is composed of $P$ blocks. Within each block, we first inject the temporal dynamic context from the Historical Actions $\tilde{\textbf{A}}$ into the Denoising Targets N through Causal Temporal Attention (CTA), generating intermediate features $\mathbf{N}_{t}$ . Following this, Visual-Action Cross Attention (VACA) is utilized to incorporate spatial constraints from the degraded observation O into these $\mathbf{N}_{t}$, generating intermediate features $\mathbf{N}_{ts}$. These $\mathbf{N}_{ts}$ are then further processed by a Multi-layer Perceptron (MLP). It is worth noting that we place the denoising timestep embedding $t$ in the VACA layer, rather than the CTA layer in order to ensure the cached features of the Historical Actions $\tilde{\textbf{A}}$ can be correctly shared during the whole denoising process (Sec. 3.2). The entire process within each block can be described as follows,
@@
where $\mathbf{N}_{o}$ represents the output features of this block. LN denotes the Layer Normalization operation, and Enc denotes the 2D or 3D data encoder.
@@ 3.2 Chunk-wise Autoregressive Inference with Cache Sharing
$\mathbf{A}^{k}\sim p_{\theta}(\mathbf{A}^{k}|\mathbf{N}^{k},\tilde{\textbf{A}}^{k})$ (6)
@@
Cache Sharing. In the AR-k step, the Historical Actions $\tilde{\textbf{A}}^{k}$ of length $L$ can be divided into two parts: Cached Historical Actions $\tilde{\textbf{A}}^{k}_{0:l-1}$ and Uncached Historical Actions $\tilde{\textbf{A}}^{k}_{l:L-1}$, based on whether their Key and Value representations have been cached up to this step. Here, $l$ denotes the length of the Cached Historical Actions. During denoising, the Key and Value representations for the Cached Historical Actions $\tilde{\textbf{A}}^{k}_{0:l-1}$ have been precomputed in prior autoregressive steps and can be directly used in attention computation. We denote them as $\mathbf{K}_{0:l-1}^{(0)}$ and $\mathbf{V}_{0:l-1}^{(0)}$, where $(\cdot)$ indicates the denoising timestep and $(0)$ means these Key and Value representations correspond to the final Target Actions generated in prior steps. For the Uncached Historical Actions $\tilde{\textbf{A}}^{k}_{l:L-1}$ and Denoising Targets $\textbf{N}^{k}$, we utilize a QKV_MLP layer to extract their Query, Key, and Value representations,
$\{\mathbf{Q}_{l:L-1}^{(0)},\mathbf{K}_{l:L-1}^{(0)},\mathbf{V}_{l:L-1}^{(0)}\}$ $=\texttt{QKV\_MLP}~(\tilde{\textbf{A}}^{k}_{l:L-1})$ (7)
$\{\mathbf{Q}_{L:L+M-1}^{(t)},\mathbf{K}_{L:L+M-1}^{(t)},\mathbf{V}_{L:L+M-1}^{(t)}\}$ $=\texttt{QKV\_MLP}~(\textbf{N}^{k})$ (8)
@@ 4 Simulation Experiments
4.1 Quantitative Results
@@ 4.2 Ablation Study
Figure 4: Visualization of the Effectiveness and Efficiency of Our CDP. (a) The performance of DP3 collapses as observation degrades, whereas our CDP remains nearly unaffected. (b) Our Cache Sharing Mechanism yields inference-time speed-ups that scale approximately linearly with the length of the Historical Actions.
In this section, we conduct experiments on the Lift Barrier task from RoboFactory [46] to rigorously evaluate both the effectiveness and efficiency of our proposed CDP.
Robustness to Degraded Observation Data. In this experiment, we introduce varying levels of noise to the low-resolution point clouds (size: $64\times 3$), and assess the success rates of the compared methods when provided with these noisy inputs. Fig. 4 demonstrates that CDP exhibits markedly superior robustness to observation degradation. As the noise level increases, the success rate of DP3 declines precipitously, whereas our method maintains consistently high performance. This resilience is attributed to the model’s ability to leverage historical action sequences, thereby compensating for the diminished spatial constraints inherent in degraded observations.
Efficiency of Cache Sharing Mechanism. We evaluate the efficiency of our Cache Sharing Mechanism on a single RTX 4090D by measuring per-AR-step inference latency while varying the length of the Historical Data Pairs (i.e. Historical Actions $\tilde{\textbf{A}}$ and Observations O). As shown in Fig. 4, enabling Cache Sharing Mechanism consistently reduces latency, and this acceleration becomes more pronounced as the Historical Data Pairs lengthens.
@@ 6 Conclusion
Our proposed CDP effectively counteracts the impact of observation-quality degradation—arising from sensor noise, occlusions, and hardware limitations—on reliable robotic manipulation. Built upon a causal-transformer diffusion framework, the method captures critical temporal dependencies, thereby compensating for the spatial constraints diminished by degraded observations and preserving task performance. Extensive evaluations demonstrate that our CDP consistently outperforms existing approaches, attesting to its robustness and effectiveness. To facilitate real-time inference, we further introduce a Cache Sharing Mechanism that reuses pre-computed key–value attention tensors, yielding substantial reductions in the computational burden of autoregressive action generation without compromising accuracy.
Limitation. This study has not addressed tasks demanding extremely long horizons. The extension of our method to scenarios requiring sustained and coherent planning over extended temporal spans remains an avenue for future inquiry.
@@ Appendix A Simulation Experiments Details
A.1 Experiment Setup
Evaluation Metric. Following the established metrics [27], each experiment was conducted across three independent trials, utilizing seed values of 0, 1, and 2, in the Adroit, DexArt, and MetaWorld benchmarks. For each seed, the policy was evaluated over 20 episodes every 200 training epochs, and the mean of the top 5 success rates was computed. The final reported performance consists of the mean and standard deviation of these success rates across above three seeds. In the Robofactory benchmark, each experiment was executed using only one seed value (i.e. 0), and was evaluated over 100 episodes at the 300th training epoch. This comprehensive evaluation metrics ensures a fair comparison between our CDP and baseline methods, providing a clear assessment of the performance improvements achieved by our approach.
A.2 Settings
In this part, we provide detailed descriptions of all settings in our simulation experiments, including those related to the training strategy, as well as the inputs (Observations O and Historical Actions $\tilde{\textbf{A}}$) and outputs (Target Actions A).
Table 3: All Settings in Simulation Experiments. $\mathcal{N}(0,1)$ denotes Gaussian noise characterized by a mean of 0 and a variance of 1. Following [7], the Target Actions A is composed of Valid Parts and Redundant Parts.
@@ A.2 Settings
Size of Point Clouds (512, 3)
@@
Specifically, we made the following changes in the simulation experiments conducted on the RoboFactory benchmark,
  • In the 2D scenario, images were resized to (64, 64, 3) instead of (84, 84, 3). And in the 3D scenario, the size of point clouds was reduced to (128, 3), deviating from the standard configuration.
@@
  • The training epoch is set to 300 in RoboFactory benchmark.
A.3 Ablation Study
In this section, we conduct experiments on the Lift Barrier task from RoboFactory [46] to systematically examine the impact of individual hyperparameters on the performance of our CDP.
Noise Scale in Perturbed Historical Actions. To justify our design choice on noise scale in Perturbed Historical Actions, we list the quantitative results og our CDP on different noise scales in Table 4. A small noise scale reduces robustness to accumulating prediction errors, whereas an excessive magnitude distorts the underlying structure of the Historical Actions and impedes convergence. Empirically, an intermediate value of 1/6 strikes an effective balance, yielding the highest success rate across all tasks.
Table 4: Quantitative results of our proposed method on different noise scales injected in Perturbed Historical Actions.
Method Noise Scale
1/9 1/6 1/3 1 3 6
CDP3 30% 40% 36% 25% 25% 23%
Chunk Size. To justify our design choice on chunk size, we list the quantitative results of our CDP on different chunk sizes in Table 5. For simpler tasks, we recommend a Chunk Size of 8, and a larger size for more complex tasks.
Table 5: Quantitative results of our proposed method on different chunk sizes.
Method Chunk Size
1 2 4 8
CDP3 1% 5% 35% 40%
Length of the Historical Actions. The first two columns of Table 6 presents the success rates of the model for different configurations of Historical Actions length. From this table, we find that a longer sequence of Historical Actions tends to be beneficial in enhancing model performance.
Table 6: Quantitative results of our proposed methods across different length of Historical Actions. The format $_{1} -$ $_{2} -$ $_{3}$” is used to denote the following: $_{1}$ represents the length of Historical Actions. $_{2}$ and $_{3}$ represent the length of valid and redundant parts of Target Actions, respectively.
Method ? - 4 - 4 ? - 4 - 8 8 - ? - 8
4 8 16 4 8 16 2 4 8
CDP3 0% 0% 36% 0% 6% 37% 0% 6% 33%
CDP 0% 0% 37% 0 0% 39% 0% 0% 38%
Length of the Denoising Targets. The last column in Table 6 illustrates that longer Denoising Targets will provide higher gains for the model. For the whole table, we observe that the optimal length of Denoising Targets should be determined in conjunction with the length of Historical Actions. Only well-matched combinations of these action lengths can effectively enhance model performance.
A.4 Qualitative Results of Simulation Experiments
@@
Figure 5: Visualization of real-world and simulation experimental results, including Stacking Cubes, Collecting Objects, Push T, Lift Barrier, Dexart Laptop, Metaworld Box Close and Adroit Door task from the top to the bottom.
@@ B.1 Workspace
As illustrated in Fig. 6, our real-world experiments were conducted using a RealMan robotic arm fitted with a DH AG95 gripper. A stationary top-down Intel RealSense D435i RGB-D camera was employed to provide a global RGB image of the workspace. All hardware components were connected to a workstation equipped with an NVIDIA 3090 Ti GPU. This workspace enabled efficient data acquisition and evaluation of the robotic system’s performance.
@@
B.2 Experiment Setup
Demonstrations. The demonstrations utilized in our real-world experiments were meticulously collected via human teleoperation, facilitated by advanced vision-based retargeting techniques. For each task, we collected a total of 50 demonstrations, each carefully selected to encapsulate the fundamental skills and critical interactions necessary for successful task completion. This curation keeps the dataset compact while ensuring it faithfully represents the inherent complexities and challenges of each task.
Settings. All settings in our real-world experiments are list in Table 7. Similar with our simulation experiments, these settings involve the training strategy, as well as the inputs (Observations O and Historical Actions $\tilde{\textbf{A}}$) and outputs (Target Actions A).
Table 7: All Settings in Real-World Experiments.